When heated at high temperatures, a diatomic vapor disassociates as follows:

A2(g) ⇌ 2A(g)


In one experiment, a chemist finds that when 0.0490 mole of A2 was placed in a flask of volume 0.478 L at 584 K, the degree of dissociation (that is, the fraction of A2 dissociated) was 0.0246. Calculate Kc and KP for the reaction at this temperature.

Kc= 0.000252

I NEED THE ANSWER TO Kp!!!!!!!!!!!
Answer #1

Determine the relation between Kc and Kp for the equilibrium reaction as below.

Kc = [A]2/[A2]

Where the square braces denote molar concentrations.

The partial pressures for the two gaseous components can be written as

PA2V = nA2RT

Where V is the volume of the container and T is the absolute temperature of the reaction; nA2 is the number of moles of A2 in the reaction container.

Therefore,

PA2 = (nA2/V)*RT

= [A2]RT

Where [A2] = nA2/V is the molar concentration of A2 in the reaction flask.

Similarly, PA = [A]RT

The equilibrium constant Kp can be written as

Kp = (PA)2/(PA2)

= {[A]RT}2/{[A2]RT}

= [A]2/[A2]*RT

= Kc*(RT)

The value of Kc is already known. Use the known values of Kc, R and T (all without units since equilibrium constants are dimensionless) to determine Kp as

Kp = (0.000252)*(0.082 L-atm/mol.K)*(584 K)

= 0.01206

≈ 0.0121 (ans, correct to 3 sig. figs).